# Witten-O’Raifeartaigh potential revisited in the context of Warm Inflation

### v1

@@ I Introduction
But, as WI doesn’t call for a separate reheating phase, it does not face issues relating to reheating which occur in generic CI based quintessential inflation models. Quintessential inflation has previously been explored a few times in the WI setup Dimopoulos and Donaldson-Wood (2019); Rosa and Ventura (2019); Lima and Ramos (2019). In this paper too, we will explore the possibility of accommodating quintessential inflation dynamics as we have two wings of Witten-O’Raifeartaigh potential at our disposal for slow-roll dynamics. However, as we will demonstrate, in order to successfully accommodate both inflation and quintessence with this single potential, one needs to engineer a few additional features into the present model, such as (i) one needs to have two different normalizations for the two wings of the Witten-O’Raifeartaigh potential in order to account for the extreme hierarchy of energy scales between inflation and the Dark energy dominated acceleration, and (ii) facilitating the slow roll of the quintessence field in the right wing of the Witten-O’Raifeartaigh potential, too, requires additional dissipative friction, just like the inflaton field in WI in steep potentials, in order to maintain slow roll.
@@ II A Brief Discussion on Witten-O’Raifeartaigh (Cold) Inflation
Thus the slow-roll inflationary predictions were analyzed only for the right branch of this potential in Martin et al. (2014). The arrow in Fig. 1 indicates the direction along which slow roll of the inflaton proceeds in CI. Observational predictions of this model in the scalar spectral index — tensor-to-scalar ratio plane ($n_{s}$ — $r$ plane) are depicted in Fig. 156 of Martin et al. (2014).
@@ III Essentials of generic Warm Inflation dynamics
where now $\lambda$ would be the free parameter of the theory. Comparing this form of the potential with the one in Eq. (1), we see that $\phi_{0}\equiv m_{p}/\lambda$. We can see setting $\lambda\geq 1$ will probe the theories for scale $\phi_{0}$ smaller than the Planck mass, which is theoretically motivated.
With this potential, we calculate $\epsilon_{V}$, $\eta_{V}$ and $\kappa_{V}$ as
@@
We see from the above equations that when $\phi/m_{p}\gtrsim 1/(e\lambda)\sim 0.37/\lambda$, then $\epsilon_{V}$ increases and decreases when $\phi/m_{p}\lesssim 1/(e\lambda)\sim 0.37/\lambda$. However, $Q$ increases when $\phi/m_{p}\gtrsim 1/(e^{2/3}\lambda)\sim 0.51/\lambda$ and decreases for $\phi/m_{p}\lesssim 1/(e^{2/3}\lambda)\sim 0.51/\lambda$. Therefore, we need to analyze the situation in three regions:
@@ V Warm Inflation in the steep left wing of the Witten-O’Raifeartaigh potential
With the given graceful exit conditions furnished in the previous section, we see that Warm Inflation can take place in the steep left wing of the Witten-O’Raifeartaigh potential, but the last several $e$-folds of inflation will take place near the bottom of the potential as that is the region where Warm Inflation can gracefully exit. Thus, we can consider that Warm Inflation will mainly take place in the left branch but near the bottom of the potential. It is important to note that if we take values of $\lambda$ smaller than unity (which, in turn, will result in super-Planckian values of $\phi_{0}$ that are theoretically not well-supported), the steep left wing will become shallower with decreasing $\lambda$ near the bottom of the potential, and the left wing will keep becoming less steeper, as has been shown in Fig. 4. As the potential looses its steepness in its left wing with decreasing values of $\lambda$, Warm Inflation will take place in weak dissipative regime when $\lambda$ becomes lesser than unity. Thus for our model analysis we will stick to $\lambda\geq 1$ (resulting in $\phi_{0}\leq m_{p}$) which is also theoretically well-motivated and let Warm inflation take place in the strong dissipative regime.
Figure 4: The Witten-O’Raifeartaigh potential with different values of $\lambda$
@@
We have done a thorough MCMC analysis of the model given the latest cosmological data by Planck Akrami et al. (2020), ACT Louis et al. (2025); Calabrese et al. (2025) and DESI Adame et al. (2025a, b); Abdul Karim et al. (2025) using a publicly available code Cobaya. The methodology to feed in the Warm Inflationary power spectrum, given in Eq. (13), into Cobaya (or COSMOMC) by converting into a function of comoving wavenumber $k$ has been developed in Kumar and Das (2024). We followed this methodology to feed the power spectrum of our model into Cobaya and the results are furnished in Fig. 5. It can be seen from Fig. 5 that $\lambda\sim 1.029$ is preferred by the data. The other best-fit model parameters are $g_{*}\sim 6040$, $\tilde{C}_{\Upsilon}\equiv C_{\Upsilon}M^{-2}\sim 3.18\times 10^{16}\,m_{p}^{-2}$, and $V_{0}\sim 1.6\times 10^{-37}\,m_{p}^{4}$. At Hubble-crossing of the pivot scale $\phi_{*}\sim 0.57\,m_{p}$ and $Q_{*}\sim 967.9$. Warm inflation ends (i.e. when $\epsilon_{H}\sim 1$) at field value $\phi_{\rm end}\sim 0.958\,m_{p}$ (the minima of the potential for $\lambda=1.029$ lies at $\phi\simeq 0.972\,m_{p}$). These values ensure that Warm inflation is taking place in the strong dissipative regime near the bottom part of the left-wing of the potential. Inflation lasts for 42.08 $e$-folds since the Hubble-exit of the pivot scale. Our analysis yields $A_{s}\sim 2.13\times 10^{-9}$, $n_{s}=0.9762$, $\alpha_{s}=0.00162$, and $r\sim 1.6\times 10^{-30}$ (Warm Inflationary models in the strong dissipative regime, in general, produce very low tensor-to-scalar ratio). All relevant observables are in very good agreement with the latest datasets emerging from the joint analysis of Planck, ACT and DESI results. We have furnished the best-fit model parameters and the observables of our model in Table 1.
@@
Furthermore, the background analysis of the model with $\lambda\sim 1$ yield that the model smoothly transits to a radiation dominated epoch once Warm Inflation gracefully exits. This can be seen from the Fig.  6 where we have plotted the kinetic and potential energy densities of the inflaton field as well as the radiation energy density which takes over the inflaton energy density when $\epsilon_{H}$ reaches 1. It can also be seen that it takes another 0.7 $e$-folds after the graceful exit to reach a radiation dominated epoch signified by $\epsilon_{H}\sim 2$. In Fig. 7, we have shown the evolution of the $\epsilon_{H}$ parameter. We further note from the evolution of the parameter $Q$ in our model, depicted in Fig. 8, that $Q$, and therefore $\Upsilon_{r}$, steadily decreases after the universe becomes radiation dominated, and after a few $e$-foldings the scalar field will effectively cease to dissipate energy to the radiation bath. A standard radiation dominated epoch will commence thereafter.
@@ VI Scope of quintessential Dark Energy in the right wing of the Witten-O’Raifeartaigh potential
After demonstrating that Warm Inflation can successfully takes place on the steep left wing of the Witten-O’Raifeartaigh potential, we will now exploit the flatter right wing to explore the possibility of having a second accelerating phase corresponding to the present dark energy epoch. We learnt from the previous section that successful inflation on the left steep wing of the the Witten-O’Raifeartaigh potential requires $V_{0}\sim 1.6\times 10^{-37}\,m_{p}^{4}$. However dark energy ($\rho_{\rm DE}\sim 10^{-120}\,m_{p}^{4}$) is many orders of magnitude smaller than $V_{0}$. Thus, even if the inflaton field rolls up the flat right wing of the Witten-O’Raifeartaigh potential (after sourcing Warm Inflation in the left wing), the scenario still cannot mimic the present Dark Energy epoch even if the inflaton field behaves like quintessence on the right wing.
In this context it is interesting to note that in the original quintessential inflation model proposed by Peebles and Vilenkin Peebles and Vilenkin (1999), two different forms of the scalar potential were considered for inflation and quintessence:
@@
Moreover, from the analysis of the previous section, we obtained that the field value at the end of “reheating” (when $\epsilon_{H}\sim 2$) the field value of $\phi_{\rm reh}\sim 0.972\,m_{p}$, and its kinetic energy is $5.49\times 10^{-47}\,m_{p}^{4}$. With this kinetic energy the scalar field is not able to climb up the right wing of the potential sufficiently and it freezes at $\phi_{q}\sim 1.147\,m_{p}$ due to Hubble friction. Once Hubble friction diminishes significantly and the potential energy density of the quintessential field starts to dominate, the quintessence field thaws and start rolling back to the bottom of the potential along the right wing. A comparison of the terms in the Klein-Gordon equation of the quintessence field (post inflation) with the Hubble friction, depicted in Fig. 9, ensures freezing of the quintessential inflaton field for the entire post-inflationary evolution and shows that the thawing only starts relatively recently at $z\sim 100$.
@@ Acknowledgements.
The work of S.D. is supported by the Start-up Research Grant (SRG) awarded by Anusandhan National Research Foundation (ANRF), Department of Science and Technology, Government of India [File No. SRG/2023/000101/PMS]. UK warmly acknowledges the Axis Bank PhD program at Ashoka University for PhD fellowships provided by Axis Bank. SSM was supported by the STFC Consolidated Grant [ST/T000732/1] at the University of Nottingham. VS thanks the Anusandhan National Research Foundation (ANRF), India, for the National Science Chair Professorship which provided partial funding for this work. SD would like to thank Rudnei Ramos for many useful discussions over the years which have been very useful in this work. SSM gratefully acknowledges support from the IUCAA Visitor Academic Programme.

### v2

@@ I Introduction
But, as WI doesn’t call for a separate reheating phase, it does not face issues relating to reheating which occur in generic CI based quintessential inflation models. Quintessential inflation has previously been explored a few times in the WI setup Dimopoulos and Donaldson-Wood (2019); Rosa and Ventura (2019); Lima and Ramos (2019). In the first two papers Dimopoulos and Donaldson-Wood (2019); Rosa and Ventura (2019), the authors considered two separate potentials of the quintessential field for the early- and late-time evolution, just like in standard quintessential inflation models. Moreover, dissipation of the scalar field, which is a signature of WI, was considered to be effective only during the inflationary period in these models, whereas the quintessence dynamics was treated in the standard non-dissipative fashion. Later in Lima and Ramos (2019), warm quintessential inflation was achieved with a single scalar potential of the form $V(\phi)=V_{0}\exp(-\alpha\phi^{n}/m_{p}^{n})\,(n\geq 2),$ (such a potential was first explored in the context of cold quintessential inflation in Geng et al. (2015)). Moreover this paper suggested that the scalar field dissipates its energy into both radiation and matter degrees of freedom. It dissipates more to radiation during inflation (yielding WI) which ceases once inflation ends. Thereafter the scalar field continues to dissipate to matter during the later stages of evolution. Such a construction not only explains inflation and dark energy, but also produces dark matter through dissipation of the scalar field. Thus, this model attempts to explain inflation, dark matter and dark energy within one single framework. The early and late-time stability of this unified model was analyzed in Das et al. (2023). Baryogenesis was also explored in a similar warm quintessential inflation model in Basak et al. (2022). It is worth pointing out that in all these models of warm quintessential inflation, very flat potentials were chosen for the dynamics of the inflaton and the quintessence field, allowing slow-roll dynamics to take place.
In this paper too, we will explore the possibility of accommodating quintessential inflation dynamics as we have two wings of Witten-O’Raifeartaigh potential at our disposal for slow-roll dynamics. However, as we will demonstrate, in order to successfully accommodate both inflation and quintessence with this single potential, one needs to engineer a few additional features into the present model, such as (i) one needs to have two different normalizations for the two wings of the Witten-O’Raifeartaigh potential in order to account for the extreme hierarchy of energy scales between inflation and the Dark energy dominated acceleration, and (ii) facilitating the slow roll of the quintessence field in the right wing of the Witten-O’Raifeartaigh potential, too, requires additional dissipative friction, just like the inflaton field in WI in steep potentials, in order to maintain slow roll.
@@ II A Brief Discussion on Witten-O’Raifeartaigh (Cold) Inflation
Thus the slow-roll inflationary predictions were analyzed only for the right branch of this potential in Martin et al. (2014). The arrow in Fig. 1 indicates the direction along which slow roll of the inflaton proceeds in CI. Observational predictions of this model in the scalar spectral index — tensor-to-scalar ratio plane ($n_{s}$ — $r$ plane) are depicted in Fig. 156 of Martin et al. (2014). Witten-O’Raifeartaigh potential has also been explored in CI with non-minimally coupled inflaton field Santos and Silva (2022); Santos et al. (2023). However, here too, the inflation takes place in the flatter right wing of the potential.
@@ III Essentials of generic Warm Inflation dynamics
where now $\lambda$ would be the free parameter of the theory. Comparing this form of the potential with the one in Eq. (1), we see that $\phi_{0}\equiv m_{p}/\lambda$. We can see setting $\lambda\geq 1$ will probe the theories for scale $\phi_{0}$ smaller than the Planck mass, which is well motivated from an effective field theory prospective.
With this potential, we calculate $\epsilon_{V}$, $\eta_{V}$ and $\kappa_{V}$ to be
@@
We see from the above equations that when $\phi/m_{p}\gtrsim 1/(e\lambda)\sim 0.37/\lambda$, $\epsilon_{V}$ increases, while it decreases when $\phi/m_{p}\lesssim 1/(e\lambda)\sim 0.37/\lambda$. However, $Q$ increases when $\phi/m_{p}\gtrsim 1/(e^{2/3}\lambda)\sim 0.51/\lambda$ and decreases for $\phi/m_{p}\lesssim 1/(e^{2/3}\lambda)\sim 0.51/\lambda$. Therefore, we need to analyze the situation in three regions:
@@ V Warm Inflation in the steep left wing of the Witten-O’Raifeartaigh potential
With the given graceful exit conditions furnished in the previous section, we see that Warm Inflation can take place in the steep left wing of the Witten-O’Raifeartaigh potential, but the last several $e$-folds of inflation will take place near the bottom of the potential as that is the region where Warm Inflation can gracefully exit. Thus, we can consider that Warm Inflation will mainly take place in the left branch but near the bottom of the potential. It is important to note that if we take values of $\lambda$ smaller than unity (which, in turn, will result in super-Planckian values of $\phi_{0}$ that might be problematic from an effective field theory point of view), the steep left wing will become shallower with decreasing $\lambda$ near the bottom of the potential, and the left wing will keep becoming less steeper, as has been shown in Fig. 4. As the potential looses its steepness in its left wing with decreasing values of $\lambda$, Warm Inflation will take place in weak dissipative regime when $\lambda$ becomes lesser than unity. Therefore, in our following analysis, we will consider $\lambda\geq 1$ for which Warm Inflation takes place in the strong dissipative regime. Additionally, $\lambda\geq 1$ results in $\phi_{0}\leq m_{p}$, which is well justified from the effective field theory prospective.
Figure 4: The Witten-O’Raifeartaigh potential for different values of $\lambda$.
@@
We have carried out a thorough MCMC analysis of the model by combining the latest cosmological datasets from Planck Akrami et al. (2020), ACT Louis et al. (2025); Calabrese et al. (2025) and DESI Adame et al. (2025a, b); Abdul Karim et al. (2025) using the publicly available code Cobaya. The methodology to incorporate the Warm Inflationary power spectrum, given in Eq. (13), into Cobaya (or COSMOMC) by converting the power spectrum into a function of comoving wavenumber $k$ has been developed in Kumar and Das (2024). We followed this methodology to feed the power spectrum of our model into Cobaya and the results are furnished in Fig. 5. It can be seen from Fig. 5 that $\lambda\sim 1.029$ is preferred by the data. The other best-fit model parameters are $g_{*}\sim 6040$, $\tilde{C}_{\Upsilon}\equiv C_{\Upsilon}M^{-2}\sim 3.18\times 10^{16}\,m_{p}^{-2}$, and $V_{0}\sim 1.6\times 10^{-37}\,m_{p}^{4}$. At Hubble-crossing of the pivot scale $\phi_{*}\sim 0.57\,m_{p}$ and $Q_{*}\sim 967.9$. Warm inflation ends (i.e. when $\epsilon_{H}\sim 1$) at field value $\phi_{\rm end}\sim 0.958\,m_{p}$ (the minima of the potential for $\lambda=1.029$ lies at $\phi\simeq 0.972\,m_{p}$). These values ensure that Warm inflation is taking place in the strong dissipative regime near the bottom part of the left-wing of the potential. Inflation lasts for 42.08 $e$-folds since the Hubble-exit of the pivot scale. Our analysis yields $A_{s}\sim 2.13\times 10^{-9}$, $n_{s}=0.9762$, $\alpha_{s}=0.00162$, and $r\sim 1.6\times 10^{-30}$ (Warm Inflationary models in the strong dissipative regime, in general, produce very low tensor-to-scalar ratio). All relevant observables are in very good agreement with the latest datasets emerging from the joint analysis of Planck, ACT and DESI results. We have furnished the best-fit model parameters and the observables of our model in Table 1.
@@
Furthermore, the background analysis of the model with $\lambda\sim 1$ yields that the model smoothly transits to a radiation dominated epoch once Warm Inflation gracefully exits. This can be seen from the Fig.  6 where we have plotted the kinetic and potential energy densities of the inflaton field as well as the radiation energy density which takes over the inflaton energy density when $\epsilon_{H}$ reaches 1. It can also be seen that it takes another 0.7 $e$-folds after the graceful exit to reach a radiation dominated epoch signified by $\epsilon_{H}\sim 2$. In Fig. 7, we have shown the evolution of the $\epsilon_{H}$ parameter. We further note from the evolution of the parameter $Q$ in our model, depicted in Fig. 8, that $Q$, and therefore $\Upsilon_{r}$, steadily decreases after the universe becomes radiation dominated, and after a few $e$-foldings the scalar field will effectively cease to dissipate energy to the radiation bath. A standard radiation dominated epoch will commence thereafter.
@@ VI Scope of quintessential Dark Energy in the right wing of the Witten-O’Raifeartaigh potential
After demonstrating that Warm Inflation can successfully takes place on the steep left wing of the Witten-O’Raifeartaigh potential, we will now exploit the flatter right wing to explore the possibility of realizing a second accelerating phase corresponding to the present dark energy epoch. We learnt from the previous section that successful inflation on the left steep wing of the the Witten-O’Raifeartaigh potential requires $V_{0}\sim 1.6\times 10^{-37}\,m_{p}^{4}$. However dark energy ($\rho_{\rm DE}\sim 10^{-120}\,m_{p}^{4}$) is many orders of magnitude smaller than $V_{0}$. Thus, even if the inflaton field rolls up the flat right wing of the Witten-O’Raifeartaigh potential (after sourcing Warm Inflation in the left wing), the scenario still cannot mimic the present Dark Energy epoch even if the inflaton field behaves like quintessence on the right wing.
In this context it is interesting to note that in the original quintessential inflation model proposed by Peebles and Vilenkin Peebles and Vilenkin (1999), two different forms of the scalar potential were considered for the inflaton and quintessence:
@@
Moreover, from the analysis of the previous section, we obtained that the field value at the end of “reheating” (when $\epsilon_{H}\sim 2$) the field value of $\phi_{\rm reh}\sim 0.972\,m_{p}$, and its kinetic energy is $5.49\times 10^{-47}\,m_{p}^{4}$. With this kinetic energy the scalar field is not able to climb up the right wing of the potential sufficiently and it freezes at $\phi_{q}\sim 1.147\,m_{p}$ due to Hubble friction. Once Hubble friction diminishes significantly and the potential energy density of the quintessential field starts to dominate, the quintessence field thaws and begins rolling back towards the bottom of the potential along the right wing. A comparison of the terms in the Klein-Gordon equation of the quintessence field (post inflation) with the Hubble friction, depicted in Fig. 9, ensures freezing of the quintessential inflaton field for the entire post-inflationary evolution and shows that the thawing only starts relatively recently at $z\sim 100$.
@@ Acknowledgements.
The work of S.D. is supported by the Start-up Research Grant (SRG) awarded by Anusandhan National Research Foundation (ANRF), Department of Science and Technology, Government of India [File No. SRG/2023/000101/PMS]. UK warmly acknowledges the Axis Bank PhD program at Ashoka University for PhD fellowships provided by Axis Bank. UK also acknowledges the HPC facility at Ashoka University (Chanakya@Ashoka) for providing computational resources essential to this work. Additionally, UK is grateful to Kandaswamy Subramanian for insightful discussions that have contributed significantly to the development of this work. SSM is supported by IBS under the project code, IBS-R018-D3. SSM was also supported by the STFC Consolidated Grant [ST/T000732/1] at the University of Nottingham. SSM gratefully acknowledges support from the IUCAA Visitor Academic Programme. VS thanks the Anusandhan National Research Foundation (ANRF), India, for the National Science Chair Professorship which provided partial funding for this work. SD would like to thank Rudnei Ramos for many useful discussions over the years which have been very useful in this work.